OpenStudy (anonymous):

simplify: 3x-12/x squared-4x

jhonyy9 (jhonyy9):

factoriz out what you can in numerator and in denominator and will see what you will can simplifie than after this

OpenStudy (anonymous):

show me step by step please?

jhonyy9 (jhonyy9):

ok 3x -12 3x -4*3 ------- = ---------- = x^2 -4x x*x -4x how do you think what will can factorizing out in numerator and in denominator easy ?

OpenStudy (anonymous):

its a factoring problem..each rational expression has to be simplified

jhonyy9 (jhonyy9):

so than will be 3(x-4) ------ = x(x-4) so now just simplifie it and will get the right result

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so than good luck bye

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

was my pleasure yw bye